# How to Purchase Labels? A Cost-effective Approach Using Active Learning Markets

### v2

@@ Abstract
We introduce and analyse active learning markets as a way to purchase labels, in situations where analysts aim to acquire additional data to improve model fitting, or to better train models for predictive analytics applications. This comes in contrast to the many proposals that already exist to purchase features and examples. By originally formalising the market clearing as an optimisation problem, we integrate budget constraints and improvement thresholds into the label acquisition process. We focus on a single-buyer-multiple-seller setup and propose the use of two active learning strategies (variance based and query-by-committee based), paired with distinct pricing mechanisms. They are compared to benchmark baselines including random sampling and a greedy knapsack heuristic. The proposed strategies are validated on real-world datasets from two critical application domains: real estate pricing and energy forecasting. Results demonstrate the robustness of our approach, consistently achieving superior performance with fewer labels acquired compared to conventional methods. Our proposal comprises an easy-to-implement practical solution for optimising data acquisition in resource-constrained environments.
@@ 1 Introduction
In today’s data-driven economy, data is often referred to as the “new oil”, reflecting its critical role in driving technological innovation and economic growth. The global market for data analytics is projected to expand significantly in the coming years, growing from USD 348.21 billion in 2024 to USD 924.39 billion in 2032, with a compound annual growth rate of 13.9% (Fortune Business Insights, 2023). The rapid proliferation of data, generated on massive scales daily, has paved the way for advances in big data analytics, offering innovative solutions in areas such as transportation, energy, real estate, and healthcare, among others. As an example, in the UK, Citymapper leverages freely available real-time transportation data to enhance urban navigation and improve public transport efficiency (Tavmen, 2020). However, in many cases, the ownership of data is distributed among heterogeneous agents, while data is deemed a commodity that must be strategically acquired. For instance, as businesses increasingly recognize the importance of predictive capabilities, many face the challenge of acquiring the necessary volume and quality of data to train machine learning models. It is of utmost importance to develop appropriate platforms and market ecosystems supporting the sharing, valuation, and acquisition of data.
@@
In a general setup, we place ourselves in a supervised learning framework, where data for input features (equivalently referred to as explanatory variables) may be abundant, while the availability of labels (i.e., values for the target variable) is limited. To develop the idea and a proof of concept, we specifically use linear regression. Linear regression has many practical advantages, including simplicity, interpretability, and the availability of a closed-form solution for parameter estimates. Furthermore, linear regression has also been used by others as a basis to introduce various types of data markets (Dekel et al., 2010; Cummings et al., 2015a), which produces important market properties such as incentive compatibility and truthfulness. However, one should be able to broaden the concepts presented to a class of models that is more general than linear regression only. For instance, as discussed by Pinson et al. (2022), data markets developed for the linear regression case can be readily generalized to the case of regression models that are linear in their parameters; this includes, for instance, polynomial- and spline-based regression. In our framework, an analyst seeks to acquire additional labels to improve model fitting, or to improve the predictive performance of the model at hand. These may be highly dependent on the availability and quality of the labels that are used. Labels may be difficult or costly to obtain.
@@
Existing approaches to data and information markets fall into various categories. Given an analytics task at hand, one may purchase solutions for that task (as in prediction markets), or one may purchase actual data to be used as input to solving that analytics task. For descriptive and predictive analytics, this may translate to having (i) observation markets and (ii) feature markets (also referred to as regression markets). In the first case, the buyer aims to acquire complete rows (with values for both input features and the target variable) to have more data to train their model. In the second case, emphasis is placed on purchasing additional features (i.e., complete columns) to augment the model at hand and to improve their ability to explain variations in the target variable. These markets have been extensively studied (Agarwal et al., 2019; Pinson et al., 2022) and are mainly based on Shapley values to define revenues and payments. Alternatively, Han et al. (2022) proposed the use of Lasso-based estimation to elicit the willingness to pay of data buyers. Moreover, existing research often emphasizes bulk data acquisition or continuous data delivery, such as in Data as a Commodity (e.g., Amazon Mechanical Turk) or Data as a Service (e.g., Google BigQuery). These approaches fall short of addressing the need for targeted and cost-effective data acquisition, especially for machine learning applications where high-quality, informative data points are more valuable than large quantities of generic data. At the same time, the growing presence of platforms such as AWS Data Exchange, Ocean Protocol, and crowdsourcing markets like Amazon Mechanical Turk demonstrates that market-style procurement of data is already a practical reality. What remains unexplored is how to integrate active learning principles into these emerging markets to guide selective, budget-constrained acquisition of the most informative labels—which is the gap this paper addresses.
@@
In contrast, here, we introduce a new type of data market to specifically purchase labels. We refer to such markets as active learning markets, since active learning serves as a basis for acquiring labels in a value-oriented manner. In active learning markets, buyers possess full feature sets, but only a subset of values for the target variable of interest. This calls for strategic acquisition of additional labels to improve model performance. This setup closely mirrors real-world scenarios faced by companies where the resources for data acquisition (for the target variable) may be limited. Despite advances in data monetization and pricing models (Agarwal et al., 2019; Mehta et al., 2021), frameworks often overlook budget constraints and fail to propose efficient strategies for buyers to prioritize high-value data points. Furthermore, although studies such as Cummings et al. (2015b) propose variance-constrained mechanisms for data selection, they do not accommodate budget constraints and strategic selection of the data to be purchased. This gap is particularly pronounced in industries such as energy forecasting, where data acquisition is expensive, but essential to achieve high prediction accuracy (Wang et al., 2023; Settles, 2011).
@@
We bridge this research gap by introducing and analyzing active learning markets. Our core contribution is then to incorporate both overall budget constraints and minimum improvement requirements in data acquisition problems. This dual consideration is a departure from traditional data purchase problems. Second, we extend the use of the active learning principle within a data market context. Eventually, we showcased the interest of our proposal based on case studies inspired by real-world applications. To realize these contributions, we first define the market setup, introducing the market participants and the key components of the underlying resource allocation problem. We then formalize that resource allocation problem within an optimization framework. Rather than solving this optimization problem directly, we use active learning as a greedy approach to data acquisition. This is based on two alternative active learning strategies (variance-based active learning, VBAL, and query-by-committee-based active learning, QBCAL), each paired with distinct data pricing mechanisms. For comparison, we use a random sampling corrected strategy as a benchmark, and also include a greedy knapsack heuristic baseline (GK) when seller costs are heterogeneous. To validate our proposed strategies, we apply them to real-world datasets in two critical domains: real estate pricing and energy forecasting. Through these applications, we analyse and discuss the salient features of active learning markets, with the aim of underlining their practical value in optimizing data acquisition under various constraints.
@@ 2.1 Data valuation
Data, as an asset, has three unique characteristics that distinguish it from traditional goods, each presenting its own set of challenges. First, data can be easily replicated at zero cost, introducing new issues in data pricing (Acemoglu et al., 2022; Mehta et al., 2021; Liang et al., 2018; Pei, 2020; Yu and Zhang, 2017; Li et al., 2014). Second, data is generated in vast volumes and at high velocity. For instance, according to Statista (2024), the data generated in 2025 is estimated at reaching 181 zettabytes, increased by more than 150% compared to the data generated in 2023—120 zettabytes. This massive scale creates significant challenges for big data analytics (Maheshwari et al., 2021; Mariani and Wamba, 2020; Wang and Wang, 2020). Third, data ownership is distributed by companies, each with heterogeneous preferences due to differing privacy concerns and competitive dynamics (Abbas et al., 2021; Busch-Casler and Radic, 2022; OECD, 2018).
@@ 2.2 Alternative data market setups
In the first case, which we refer to as observation market, the analyst may already own, or have access to, a number of observations for both explanatory variables and the response variable they are interested in. This situation is illustrated in Figure 1 (a). There, the analyst already has the first $K$ rows and may be interested in purchasing additional rows (from the NKN-K they do not have already). In the second case, which we refer to as feature market (the term regression market has also been used in the literature, see Pinson et al. (2022), for instance), the analyst has access to all $N$ observations for $M$ input features, as well as for the response variable of interest. The analyst may then be interested in purchasing additional columns, i.e., data for additional PMP-M potential features that they do not already have access to.
@@ 2.3 Active learning market: Participants
The data analyst $A$ works with several defined components. Note that we use the term data analyst for the data buyer, to be consistent with the terminology used by, e.g., Cummings et al. (2015b). Identified data $D_{\text{L}}$ comprises the data set owned by the analyst, including both features and their corresponding target values (that is, the corresponding labels). The unlabelled data $D_{\text{U}}$ consists of data points that the analyst can access, covering values for all relevant features, but without associated labels for the target variable of interest. The analyst is willing to pay (WTP) $\phi$, which represents what the analyst is ready to spend (per unit cost, e.g., £, \$, ¥) for a unit improvement in some model performance metric $L$. For instance, both the variance of model estimates and the mean square error of model-based forecasts will be used in the following. Here, we treat $\phi$ as a constant, reflecting ex-ante scenarios in which the analyst has no prior information to differentiate among candidate labels and therefore sets a uniform value for performance improvement. This simplification follows the tradition of single-parameter environments in mechanism design, where each participant’s valuation or cost is summarized by a single scalar parameter (e.g., Ghosh and Roth (2011); Cummings et al. (2015a); Agarwal et al. (2019)). Such models provide tractability while still capturing key buyer–seller interactions. In practice, this corresponds to organizations that translate forecast errors into well-defined financial costs (e.g., inventory misallocation or excess capacity), which naturally yield a constant willingness to pay per unit of accuracy. Additionally, our analysis is conducted in the batch active learning setting; therefore, the buyer’s willingness to pay (WTP) is assumed constant over time rather than dynamically updated. The potential variation of WTP in online or adaptive settings is discussed in Section 5.2. The analyst also establishes a model performance threshold $\alpha$, which defines the maximum expected decrease in the model performance metric (assuming that the metric is negatively oriented) for which they are ready to pay. In parallel, the data purchase budget $B$ specifies the maximum amount the analyst may be able to spend to purchase data.
@@ 2.4 How the market operates
Regarding the pricing of the labels actually purchased, we consider two alternatives which are to be seen as the most extreme cases. That is, the pricing can be based either on the buyer or on the offer from the seller. In a single-buyer-multiple seller setup, one cannot rely on supply-demand equilibrium concepts, as would be the case for more general resource allocation problems with multiple buyers and multiple sellers. Here, in a seller-centric pricing approach (SC), the price is defined as $p_{j}=\eta_{j}$, which corresponds to the willingness to sell for the label $y_{j}$. Following a buyer-centric (BC) pricing approach, instead, each purchased label $y_{j}$ is priced as $p_{j}=\phi l_{j}$, i.e., reflecting the willingness to pay $\phi$ and the perceived value of the label $l_{j}$. These two approaches can be seen as extreme cases since labels will only be purchased if $\phi l_{j}\geq\eta_{j}$, and then any price $p_{j}\in[\eta_{j},\,\phi l_{j}]$ could be acceptable to both buyer and the seller. As a by-product of market clearing, the analyst can estimate a set of parameters $\hat{\theta}(D_{\text{L}},D_{\text{U}},D^{\text{ac}}_{\text{ML}})$ based on their own data, as well as acquired labels, for their regression model. The analyst also deduces the estimate of the overall loss function by acquiring $D^{\text{ac}}_{\text{ML}}\subseteq D_{\text{ML}}$. We write $l_{j}$ the loss reduction induced by acquiring the label $y_{j}$.
@@ 2.5 Data purchase as an optimisation problem
Although the simplified structure in (5) superficially resembles a knapsack formulation, applying knapsack algorithms directly is infeasible in our setting. Knapsack methods—exact, approximate, or heuristic (Martello and Toth, 1990; Bourdache and Perny, 2019)—all require that the utilities of all items be known in advance. In our context, these utilities correspond to the realised improvement from labeling each point, which can only be observed after the label has been purchased. This makes classical knapsack approaches unsuitable for strategic and limited label acquisition. Moreover, even if approximate utilities were available, the combinatorial nature of the problem combined with the scale of modern datasets renders exact solvers and approximation schemes computationally burdensome. Nevertheless, to provide a reference baseline, we later report a greedy knapsack-style heuristic (GK) that ranks candidates using ex-ante proxies rather than realised utilities; it is therefore not a direct knapsack solution to (5) and offers no optimality guarantee. In contrast, active learning provides a scalable and adaptive alternative: it uses estimators of label utility that can be updated as labels are acquired, without requiring all labels to be observed upfront (Cacciarelli and Kulahci, 2024). Active learning trades global optimality guarantees for adaptability and computational feasibility, making it more practical for real-world large-scale label acquisition tasks.
@@
Deriving full optimality or regret guarantees for (5) is challenging because the marginal utilities $\{l_{j}\}$ are revealed only after purchase and are state-dependent (they vary with the acquired set), so standard analyses that assume a static set function do not directly apply. Nevertheless, partial theoretical insights may be obtainable. For instance, assume that (i) costs are bounded or drawn i.i.d. from a distribution independent of informativeness, i.e. $\eta_{j}\in[\eta_{\min},\eta_{\max}]$, and (ii) a pre-purchase proxy $\widehat{l}_{j}$ is bounded and (approximately) unbiased for the true marginal improvement $l_{j}$. Under such assumptions, a proxy-based cost-weighted greedy AL rule that ranks candidates by $\widehat{l}_{j}/\eta_{j}$ aligns with cost-sensitive greedy query-tree analyses that select by a benefit–cost ratio (e.g., shrinkage-cost ratio) (Guillory and Bilmes, 2009), and can be studied via (a) approximation-style guarantees against optimal (or baseline) query trees under simplified models (Dasgupta, 2004), or (b) myopic (one-step) greedy optimality properties implied by selecting the maximising ratio at each step (Guillory and Bilmes, 2009). Our market setting does not satisfy these static-utility assumptions: both proxies and realised improvements evolve as labels are acquired, and acceptance decisions couple $\eta_{j}$ and $l_{j}$ through (5c)–(5e). Establishing formal bounds would therefore require additional structural conditions (e.g., diminishing returns / adaptive submodularity) under which greedy policies are provably near-optimal (Golovin and Andreas, 2011), which we leave for future work.

### v3

@@ Abstract
We introduce and analyse active learning markets as a way to purchase labels, in situations where analysts aim to acquire additional data to improve model fitting, or to better train models for predictive analytics applications. This comes in contrast to the many proposals that already exist to purchase features and examples. By originally formalizing the market clearing as an optimisation problem, we integrate budget constraints and improvement thresholds into the label acquisition process. We focus on a single-buyer-multiple-seller setup and propose the use of two active learning strategies (variance based and query-by-committee based), paired with distinct pricing mechanisms. They are compared to benchmark baselines including random sampling and a greedy knapsack heuristic. The proposed strategies are validated on real-world datasets from two critical application domains: real estate pricing and energy forecasting. Results demonstrate the robustness of our approach, consistently achieving superior performance with fewer labels acquired compared to conventional methods. Our proposal comprises an easy-to-implement practical solution for optimizing data acquisition in resource-constrained environments.
@@ 1 Introduction
In today’s data-driven economy, data is often referred to as the “new oil”, reflecting its critical role in driving technological innovation and economic growth. The global market for data analytics is projected to expand significantly in the coming years, growing from USD 348.21 billion in 2024 to USD 924.39 billion in 2032, with a compound annual growth rate of 13.9% [17]. The rapid proliferation of data, generated on massive scales daily, has paved the way for advances in big data analytics, offering innovative solutions in areas such as transportation, energy, real estate, and healthcare, among others. As an example, in the UK, Citymapper leverages freely available real-time transportation data to enhance urban navigation and improve public transport efficiency [35]. However, in many cases, the ownership of data is distributed among heterogeneous agents, while data is deemed a commodity that must be strategically acquired. For instance, as businesses increasingly recognize the importance of predictive capabilities, many face the challenge of acquiring the necessary volume and quality of data to train machine learning models. It is of utmost importance to develop appropriate platforms and market ecosystems supporting the sharing, valuation, and acquisition of data.
@@
In a general setup, we place ourselves in a supervised learning framework, where data for input features (equivalently referred to as explanatory variables) may be abundant, while the availability of labels (i.e., values for the target variable) is limited. To develop the idea and a proof of concept, we specifically use linear regression. Linear regression has many practical advantages, including simplicity, interpretability, and the availability of a closed-form solution for parameter estimates. Furthermore, linear regression has also been used by others as a basis to introduce various types of data markets [16, 13], which produces important market properties such as incentive compatibility and truthfulness. However, one should be able to broaden the concepts presented to a class of models that is more general than linear regression only. For instance, as discussed by Pinson et al. [32], data markets developed for the linear regression case can be readily generalized to the case of regression models that are linear in their parameters; this includes, for instance, polynomial- and spline-based regression. In our framework, an analyst seeks to acquire additional labels to improve model fitting, or to improve the predictive performance of the model at hand. These may be highly dependent on the availability and quality of the labels that are used. Labels may be difficult or costly to obtain.
@@
Existing approaches to data and information markets fall into various categories. Given an analytics task at hand, one may purchase solutions for that task (as in prediction markets), or one may purchase actual data to be used as input to solving that analytics task. For descriptive and predictive analytics, this may translate to having (i) observation markets and (ii) feature markets (also referred to as regression markets). In the first case, the buyer aims to acquire complete rows (with values for both input features and the target variable) to have more data to train their model. In the second case, emphasis is placed on purchasing additional features (i.e., complete columns) to augment the model at hand and to improve their ability to explain variations in the target variable. These markets have been extensively studied [4, 32] and are mainly based on Shapley values to define revenues and payments. Alternatively, Han et al. [22] proposed the use of Lasso-based estimation to elicit the willingness to pay of data buyers. Moreover, existing research often emphasizes bulk data acquisition or continuous data delivery, such as in Data as a Commodity (e.g., Amazon Mechanical Turk) or Data as a Service (e.g., Google BigQuery). These approaches fall short of addressing the need for targeted and cost-effective data acquisition, especially for machine learning applications where high-quality, informative data points are more valuable than large quantities of generic data. At the same time, the growing presence of platforms such as AWS Data Exchange, Ocean Protocol, and crowdsourcing markets like Amazon Mechanical Turk demonstrates that market-style procurement of data is already a practical reality. What remains unexplored is how to integrate active learning principles into these emerging markets to guide selective, budget-constrained acquisition of the most informative labels—which is the gap this paper addresses.
@@
In contrast, here, we introduce a new type of data market to specifically purchase labels. We refer to such markets as active learning markets, since active learning serves as a basis for acquiring labels in a value-oriented manner. In active learning markets, buyers possess full feature sets, but only a subset of values for the target variable of interest. This calls for strategic acquisition of additional labels to improve model performance. This setup closely mirrors real-world scenarios faced by companies where the resources for data acquisition (for the target variable) may be limited. Despite advances in data monetization and pricing models [4, 29], frameworks often overlook budget constraints and fail to propose efficient strategies for buyers to prioritize high-value data points. Furthermore, although studies such as Cummings et al. [14] propose variance-constrained mechanisms for data selection, they do not accommodate budget constraints and strategic selection of the data to be purchased. This gap is particularly pronounced in industries such as energy forecasting, where data acquisition is expensive, but essential to achieve high prediction accuracy [38, 33].
@@
We bridge this research gap by introducing and analyzing active learning markets. Our core contribution is then to incorporate both overall budget constraints and minimum improvement requirements in data acquisition problems. This dual consideration is a departure from traditional data purchase problems. Second, we extend the use of the active learning principle within a data market context. Eventually, we showcased the interest of our proposal based on case studies inspired by real-world applications. To realize these contributions, we first define the market setup, introducing the market participants and the key components of the underlying resource allocation problem. We then formalize that resource allocation problem within an optimization framework. Rather than solving this optimization problem directly, we use active learning as a greedy approach to data acquisition. This is based on two alternative active learning strategies (variance-based active learning, VBAL, and query-by-committee-based active learning, QBCAL), each paired with distinct data pricing mechanisms. For comparison, we use a random sampling corrected strategy as a benchmark, and also include a greedy knapsack heuristic baseline (GK) when seller costs are heterogenous . To validate our proposed strategies, we apply them to real-world datasets in two critical domains: real estate pricing and energy forecasting. Through these applications, we analyse and discuss the salient features of active learning markets, with the aim of underlining their practical value in optimizing data acquisition under various constraints.
@@ 2.1 Data valuation
Data, as an asset, has three unique characteristics that distinguish it from traditional goods, each presenting its own set of challenges. First, data can be easily replicated at zero cost, introducing new issues in data pricing [3, 29, 24, 31, 41, 23]. Second, data is generated in vast volumes and at high velocity. For instance, according to Statista [34], the data generated in 2025 is estimated at reaching 181 zettabytes, increased by more than 150% compared to the data generated in 2023—120 zettabytes. This massive scale creates significant challenges for big data analytics [26, 27, 39]. Third, data ownership is distributed by companies, each with heterogeneous preferences due to differing privacy concerns and competitive dynamics [1, 10, 30].
@@ 2.2 Alternative data market setups
In the first case, which we refer to as observation market, the analyst may already own, or have access to, a number of observations for both explanatory variables and the response variable they are interested in. This situation is illustrated in Figure 1 (a). There, the analyst already has the first $K$ rows and may be interested in purchasing additional rows (from the NKN-K they do not have already). In the second case, which we refer to as feature market (the term regression market has also been used in the literature, see Pinson et al. [32], for instance), the analyst has access to all $N$ observations for $M$ input features, as well as for the response variable of interest. The analyst may then be interested in purchasing additional columns, i.e., data for additional PMP-M potential features that they do not already have access to.
@@ 2.3 Active learning market: Participants
The data analyst $A$ works with several defined components. Note that we use the term data analyst for the data buyer, to be consistent with the terminology used by, e.g., Cummings et al. [14]. Identified data $D_{\text{L}}$ comprises the data set owned by the analyst, including both features and their corresponding target values (that is, the corresponding labels). The unlabelled data $D_{\text{U}}$ consists of data points that the analyst can access, covering values for all relevant features, but without associated labels for the target variable of interest. The analyst is willing to pay (WTP) $\phi$, which represents what the analyst is ready to spend (per unit cost, e.g., £, \$, ¥) for a unit improvement in some model performance metric $L$. For instance, both the variance of model estimates and the mean square error of model-based forecasts will be used in the following. Here, we treat $\phi$ as a constant, reflecting ex-ante scenarios in which the analyst has no prior information to differentiate among candidate labels and therefore sets a uniform value for performance improvement. This simplification follows the tradition of single-parameter environments in mechanism design, where each participant’s valuation or cost is summarized by a single scalar parameter (e.g., Ghosh and Roth [19], Cummings et al. [13], Agarwal et al. [4]). Such models provide tractability while still capturing key buyer–seller interactions. In practice, this corresponds to organizations that translate forecast errors into well-defined financial costs (e.g., inventory misallocation or excess capacity), which naturally yield a constant willingness to pay per unit of accuracy. Additionally, our analysis is conducted in the batch active learning setting; therefore, the buyer’s willingness to pay (WTP) is assumed constant over time rather than dynamically updated. The potential variation of WTP in online or adaptive settings is discussed in Section 5.2. The analyst also establishes a model performance threshold $\alpha$, which defines the maximum expected decrease in the model performance metric (assuming that the metric is negatively oriented) for which they are ready to pay. In parallel, the data purchase budget $B$ specifies the maximum amount the analyst may be able to spend to purchase data.
@@ 2.4 How the market operates
Regarding the pricing of the labels actually purchased, we consider two alternatives which are to be seen as the most extreme cases. That is, the pricing can be based either on the buyer or on the offer from the seller. In a single-buyer-multiple seller setup, one cannot rely on supply-demand equilibrium concepts, as would be the case for more general resource allocation problems with multiple buyers and multiple sellers. Here, in a seller-centric pricing approach (SC), the price is defined as $p_{j}=\eta_{j}$, which corresponds to the willingness to sell for the label $y_{j}$. Following a buyer-centric (BC) pricing approach, instead, each purchased label $y_{j}$ is priced as $p_{j}=\phi l_{j}$, i.e., reflecting the willingness to pay $\phi$ and the perceived value of the label $l_{j}$. These two approaches can be seen as extreme cases since labels will only be purchased if $\phi l_{j}\geq\eta_{j}$, and then any price $p_{j}\in[\eta_{j},\,\phi l_{j}]$ could be acceptable to both both buyer and the seller. As a by-product of market clearing, the analyst can estimate a set of parameters $\hat{\theta}(D_{\text{L}},D_{\text{U}},D^{\text{ac}}_{\text{ML}})$ based on their own data, as well as acquired labels, for their regression model. The analyst also deduces the estimate of the overall loss function by acquiring $D^{\text{ac}}_{\text{ML}}\subseteq D_{\text{ML}}$. We write $l_{j}$ the loss reduction induced by acquiring the label $y_{j}$.
@@ 2.5 Data purchase as an optimisation problem
Although the simplified structure in (5) superficially resembles a knapsack formulation, applying knapsack algorithms directly is infeasible in our setting. Knapsack methods—exact, approximate, or heuristic [28, 9]—all require that the utilities of all items be known in advance. In our context, these utilities correspond to the realised improvement from labeling each point, which can only be observed after the label has been purchased. This makes classical knapsack approaches unsuitable for strategic and limited label acquisition. Moreover, even if approximate utilities were available, the combinatorial nature of the problem combined with the scale of modern datasets renders exact solvers and approximation schemes computationally burdensome. Nevertheless, to provide a reference baseline, we later report a greedy knapsack-style heuristic (GK) that ranks candidates using ex-ante proxies rather than realised utilities; it is therefore not a direct knapsack solution to (5) and offers no optimality guarantee. In contrast, active learning provides a scalable and adaptive alternative: it uses estimators of label utility that can be updated as labels are acquired, without requiring all labels to be observed upfront [11]. Active learning trades global optimality guarantees for adaptability and computational feasibility, making it more practical for real-world large-scale label acquisition tasks.
@@
Deriving full optimality or regret guarantees for (5) is challenging because the marginal utilities $\{l_{j}\}$ are revealed only after purchase and are state-dependent (they vary with the acquired set), so standard analyses that assume a static set function do not directly apply. Nevertheless, partial theoretical insights may be obtainable. For instance, assume that (i) costs are bounded or drawn i.i.d. from a distribution independent of informativeness, i.e. $\eta_{j}\in[\eta_{\min},\eta_{\max}]$, and (ii) a pre-purchase proxy $\widehat{l}_{j}$ is bounded and (approximately) unbiased for the true marginal improvement $l_{j}$. Under such assumptions, a proxy-based cost-weighted greedy AL rule that ranks candidates by $\widehat{l}_{j}/\eta_{j}$ aligns with cost-sensitive greedy query-tree analyses that select by a benefit–cost ratio (e.g., shrinkage-cost ratio) [21], and can be studied via (a) approximation-style guarantees against optimal (or baseline) query trees under simplified models [15], or (b) myopic (one-step) greedy optimality properties implied by selecting the maximising ratio at each step [21]. Our market setting does not satisfy these static-utility assumptions: both proxies and realised improvements evolve as labels are acquired, and acceptance decisions couple $\eta_{j}$ and $l_{j}$ through (5c)–(5e). Establishing formal bounds would therefore require additional structural conditions (e.g., diminishing returns / adaptive submodularity) under which greedy policies are provably near-optimal [20], which we leave for future work.